# AI-accelerated Discovery of Altermagnetic Materials

### v3

@@ Results
Based on the proposed AI search engine, we successfully discovered 50 new altermagnetic materials including 16 metals and 34 insulators (see Table 1). The computational results for most of the newly discovered altermagnetic materials (e.g., the first 23 materials listed in Table 1) are shown in the Supplementary Note B and Supplementary Figs. S.6–S.17. Moreover, the $d$-wave, $g$-wave, and $i$-wave altermagnets can be found in the predicted 50 altermagnetic materials as shown in Table 1. In particular, we predicted 4 $i$-wave altermagnetic materials for the first time.
The 11 metallic altermagnetic materials can be divided into two classes according to whether the integral of the Berry curvature of the occupied states over the Brillouin zone is zero, which depends on the symmetry of the altermagnetic materials. Since the easy magnetization axes of these materials are in the $x$-$y$ plane, the 8 metallic altermagnetic materials ($\text{Nb}_{2}\text{Fe}\text{B}_{2}$, $\text{Ta}_{2}\text{Fe}\text{B}_{2}$, $\text{Nd}\text{B}_{2}\text{C}_{2}$, $\text{Mg}_{2}\text{Fe}\text{Ir}_{5}\text{B}_{2}$, $\text{Mg}_{2}\text{Mn}\text{Ir}_{5}\text{B}_{2}$, $\text{Mg}_{2}\text{Ni}\text{Ir}_{5}\text{B}_{2}$, $\text{Sc}_{2}\text{V}\text{Ir}_{5}\text{B}_{2}$, $\text{Sc}_{2}\text{Mn}\text{Ir}_{5}\text{B}_{2}$) have nonzero Berry curvature for the integral of the occupied states over the Brillouin zone according to magnetic point group symmetry, implying that odd-under-time-reversal responses (e.g., anomalous Hall and Kerr effects) can be realized in these materials. Especially, the calculated intrinsic anomalous Hall conductance of altermagnet $\text{Nb}_{2}\text{Fe}\text{B}_{2}$ is $-100\Omega^{-1}cm^{-1}$ [40], which is the same order of magnitude as those of ferromagnetic metals. Since the remaining 3 altermagnetic materials ($\text{Nd}\text{B}_{2}\text{C}_{2}$, $\text{Sc}_{2}\text{Mn}\text{Ir}_{5}\text{B}_{2}$, $\text{Mg}_{2}\text{Ni}\text{Ir}_{5}\text{B}_{2}$), whose easy magnetization axes are in the $z$ direction, have zero Berry curvature for the integral of the occupied states over the Brillouin zone, the anomalous Hall effect is not observed.
@@
On the other hand, ferromagnetic semiconductors that have spintronic and transistor functionalities could be applied to the next generation of electronic devices. However, the ferromagnets usually are metals with a high Curie temperature and hold no brief for insulators with a high Curie temperature. Altermagnets with compensated antiparallel sublattices not only are in favor of insulators with a high Neel temperature but also have spintronic functionality [2]. Thus, altermagnets open a new pathway to bypass the difficulties of ferromagnets. Here, we employed the LDA+U method [41] to predict 26 altermagnetic semiconductors (see Supplementary Table S.2). Furthermore, the altermagnetic $\text{Fe}\text{H}\text{O}_{2}$ may be a spin-triplet excitonic phase. From Fig. 4i, we observe that there is large spin splitting on the $\rm{T-\Gamma-T}$ directions and the spin of valence and conduction band are opposite, which may result in the spin-triplet excitonic phase [42]. Moreover, the energy of altermagnetic state (AFM1) is much lower than that of the other three magnetic states (see Fig. 4g), indicating that $\text{Fe}\text{H}\text{O}_{2}$ may have a Neel temperature above the room temperature. Thus, the altermagnetic $\text{Fe}\text{H}\text{O}_{2}$ is a very interesting material that, we believe, will attract both theoretical and experimental interests. In addition, although $\text{Nd}\text{Ru}\text{O}_{3}$ is a altermagnetic semimetal, it is bandgap in the high-symmetry direction of spin splitting, and its valence and conduction bands have opposite spins (Supplementary Fig. S.5e and f). Thus, $\text{Nd}\text{Ru}\text{O}_{3}$ may be a Bardeen Cooper Schrieffer (BCS) type triplet exciton insulator [43]. In the following, we present in detail two altermagnetic materials which are metal and semiconductor, respectively.
@@
The other altermagnetic material that we would like to mention is $\text{Na}\text{Fe}\text{O}_{2}$. The crystal structure of $\text{Na}\text{Fe}\text{O}_{2}$ is shown in Fig. 6a-d with space group $P/4_{1}2_{1}2\leavevmode\nobreak\ (92)$ symmetry. The corresponding elementary symmetry operations are $C_{4z}(\frac{1}{2},\frac{1}{2},\frac{1}{4})$ and $C_{2x}(\frac{1}{2},\frac{1}{2},\frac{3}{4})$, which yield the point group $D_{4}$. Since the $d$ orbitals of $Fe$ are half occupied and the angle between $Fe-O-Fe$ is 136 degrees in $\text{Na}\text{Fe}\text{O}_{2}$, the superexchange interactions result in the nearest neighbor $Fe$ ions with opposite magnetic moments and the next neighbor $Fe$ ions with the same magnetic moments. Hence, the magnetic ground state of $\text{Na}\text{Fe}\text{O}_{2}$ will be the AFM1 (see Fig. 6b). To verify our theoretical analysis, we consider four different magnetic structures, which are shown in Fig. 6a-d. It can be seen that the magnetic structure AFM1 is always in the ground state of $\text{Na}\text{Fe}\text{O}_{2}$ under different correlation interactions $U$ (see Fig. 6e). Moreover, the energy of the AFM1 state is much lower than that of the other three magnetic states (Fig. 6e), implying that $\text{Na}\text{Fe}\text{O}_{2}$ may have a Neel temperature above the room temperature. It is shown in Fig. 6b that the magnetic and crystal primitive cells of $\text{Na}\text{Fe}\text{O}_{2}$ are the same, which break $\{C_{2}^{\perp}||\tau\}$ spin symmetry. Thus, $\text{Na}\text{Fe}\text{O}_{2}$ is an altermagnetic material due to the lack of space-inversion symmetry.
@@
We also calculated the electronic band structure along the high-symmetry directions. Fig. 6f shows that the altermagnet $\text{Na}\text{Fe}\text{O}_{2}$ is a semiconductor with a band gap of 2.75 eV. The spin-degenerate bands on the $\Gamma-{\rm{X,\leavevmode\nobreak\ M}}^{*}-{\rm{A}}$ and $\rm{Z-R}$ directions (the $\rm{X-M}$ direction) are protected by the spin symmetry $\{C_{2}^{\perp}||C_{2y}(\frac{1}{2},\frac{1}{2},\frac{1}{4})\}$ ( $\{C_{2}^{\perp}||C_{2y}(\frac{1}{2},\frac{1}{2},\frac{3}{4})\}$) (see Fig. 6f). In fact, the spin symmetry $\{T||C_{2y}T(\frac{1}{2},\frac{1}{2},\frac{1}{4})\}$ ( $\{T||C_{2x}T(\frac{1}{2},\frac{1}{2},\frac{3}{4})\}$) can protect spin degeneracy of bands on the $k_{y}=0$ and $\pi$ (the $k_{x}=0$ and $\pi$) planes. That is to say, the altermagnet $\text{Na}\text{Fe}\text{O}_{2}$ has four nodal surfaces in the Brillouin Zone. Thus, $\text{Na}\text{Fe}\text{O}_{2}$ is a $d$-wave altermagnet which is reflected by the spin-splitting bands on the $\rm{M}-\Gamma-\rm{M}^{*}$ directions. Considering the $d$-wave altermagnets allowing unique spin current by electrical means [4], the altermagnet $\text{Na}\text{Fe}\text{O}_{2}$ may have both spintronic and transistor functionalities at the room temperature.
@@ Discussion
AI approaches have shown ground-breaking capabilities in the discovery of materials in a large search space. An intractable challenge faced by AI lies in the shortage of sufficient labels or positive samples, e.g., in the case of the discovery of altermagnetic materials. We herein introduced an AI search engine that combines pre-trained crystal models (GNN pre-training and optimal transport theory) and physics-based methods (symmetry analysis and first-principles electronic structure calculations) to discover new altermagnetic materials with specific properties under minimal labeled sample conditions. Among 91,649 possible candidates, we identified 50 new altermagnetic materials covering metal, semiconductor, and insulator. Meanwhile, the proposed AI search engine also has the few-shot learning ability. For example, it is capable of predicting 25 altermagnetic materials only based on 14 positive samples (see Supplementary Note C). We observed various novel physical properties in these newly discovered altermagnetic materials, e.g., anomalous Hall effect, anomalous Kerr effect, and topological property. It is noted that 4 out of these 50 altermagnetic materials are $i$-wave types, which are discovered for the first time filling a gap in the literature. We demonstrate that the AI search engine is capable of uncovering a set of altermagnetic materials with unique properties, highlighting its potential for accelerated discovery of the materials with targeting properties.
@@
There still remain some potential limitations associated with the AI search engine. Firstly, we have to admit that the issue of imbalance between positive and negative samples during the fine-tuning stage exists, primarily due to the scarcity of known positive samples. We have also discussed the potential intrinsic error and computational cost associated with this AI search engine (see Supplementary Note C). Utilizing the translational and rotational symmetries of crystals to augment positive sample data may help address this challenge, which will be demonstrated in our future work. Secondly, if the materials exhibit distinct magnetic phases at varying temperatures, such as Mn5Si3, whose low-temperature spin pattern is reported to be non-collinear and high-temperature spine pattern as antiferromagnetic  [44], our current model is unable to predict their magnetic properties. In such a case, if the material with temperature-driven magnetic transition is used as the positive sample to train the AI model, the prediction might involve possible bias. We will improve our classifier model by considering temperature as a conditional input parameter to enhance its capability in screening and predicting materials with temperature-driven properties in the future. Another limitation is that we have not yet found ideal altermagnetic topological insulators and altermagnetic topological semimetals (such as odd-under-time-reversal Dirac, and sixfold semimetals). Employing the decoder based on the pre-trained model to generate potential altermagnetic materials holds promise in overcoming this challenge. Furthermore, adopting a multimodal pre-training approach offers the potential to further enhance the accuracy of model predictions. The current pre-training only considers the single modality of crystal structure information. Leveraging information from other modalities (such as textual descriptions of crystal structures) may enhance the performance of the pre-trained model. These methods will be further explored in our future research endeavors.
@@
Compared to conventional methods, this AI search engine significantly improves the efficiency of altermagnetic materials discovery. The success of this engine lies not only in its predictive capabilities but also in its ability to leverage extensive crystal structure data and deep learning techniques, allowing for pre-training without explicit reference to underlying physical laws, to reveal complex correlations and patterns in new materials. Although over 100 altermagnetic materials have been confirmed by the symmetry analysis approach [19], the urgency of discovering a variety of new altermagnetic materials with different properties still remains. Based on these existing altermagnetic materials confirmed by the spin space group, our AI search engine could predict many more altermagnetic materials, among which we expect to find a variety of altermagnets beyond the MAGNDATA database. For example, over three hundred additional candidate materials (unconfirmed yet by DFT calculations) were predicted by AI search engine and listed in our GitHub repository [^1] https://github.com/zfgao66/MatAltMag for open research. The proposed engine is also applicable to other types of materials whose properties are strongly correlated to their crystal structures, such as the Bardeen-Cooper-Schrieffer (BCS) superconductive materials [45], ferromagnetic semi-conductor [46], and high-temperature superconductive materials [47], among others. We will demonstrate the potential of our proposed model for discovering other materials in our future study. We envision this effort may present new opportunities in the field of material discovery across different disciplines.

### v4

@@ 50 $\text{Li}\text{Fe}\text{F}_{4}$ $P2_{1}/c~{}(14)$ $d$-wave I mp-755632 NA
Based on the proposed AI search engine, we successfully discovered 50 new altermagnetic materials including 16 metals and 34 insulators (see Table 1). The computational results for most of the newly discovered altermagnetic materials (e.g., the first 23 materials listed in Table 1) are shown in the Supplementary Appendix Note B and Supplementary Appendix Figs. S.6–S.17. Moreover, the $d$-wave, $g$-wave, and $i$-wave altermagnets can be found in the predicted 50 altermagnetic materials as shown in Table 1. In particular, we predicted 4 $i$-wave altermagnetic materials for the first time.
Figure 3: The crystal and electronic structure of the altermagnets. a, The NdB2C2 crystal primitive cell with magnetic structure. b and c, The electronic band structure of altermagnetic NdB2C2. The electronic structure is calculated under correlation interaction $\rm{U=5~{}eV}$. d, The Sc2MnIr5B2 crystal primitive cell with magnetic structure. e and f, The electronic band structure of altermagnetic Sc2MnIr5B2 without and with SOC, respectively. The electronic structure is calculated under correlation interaction $\rm{U=4~{}eV}$. g, The Mg2NiIr5B2 crystal primitive cell with magnetic structure. h and i, The electronic band structure of altermagnetic Mg2NiIr5B2 without and with SOC, respectively. The electronic structure is calculated under correlation interaction $\rm{U=6.56~{}eV}$. The red and blue lines represent spin-up and spin-down energy bands, respectively.
The 16 metallic altermagnetic materials can be divided into two classes according to whether the integral of the Berry curvature of the occupied states over the Brillouin zone is zero, which depends on the symmetry of the altermagnetic materials. Since the easy magnetization axes of these materials are in the $x$-$y$ plane, the 8 metallic altermagnetic materials ($\text{Nb}_{2}\text{Fe}\text{B}_{2}$, $\text{Ta}_{2}\text{Fe}\text{B}_{2}$, $\text{Nd}\text{B}_{2}\text{C}_{2}$, $\text{Mg}_{2}\text{Fe}\text{Ir}_{5}\text{B}_{2}$, $\text{Mg}_{2}\text{Mn}\text{Ir}_{5}\text{B}_{2}$, $\text{Mg}_{2}\text{Ni}\text{Ir}_{5}\text{B}_{2}$, $\text{Sc}_{2}\text{V}\text{Ir}_{5}\text{B}_{2}$, $\text{Sc}_{2}\text{Mn}\text{Ir}_{5}\text{B}_{2}$) have nonzero Berry curvature for the integral of the occupied states over the Brillouin zone according to magnetic point group symmetry, implying that odd-under-time-reversal responses (e.g., anomalous Hall and Kerr effects) can be realized in these materials. Especially, the calculated intrinsic anomalous Hall conductance of altermagnet $\text{Nb}_{2}\text{Fe}\text{B}_{2}$ is $-100\Omega^{-1}cm^{-1}$ [83], which is the same order of magnitude as those of ferromagnetic metals. Since the 3 altermagnetic materials ($\text{Nd}\text{B}_{2}\text{C}_{2}$, $\text{Sc}_{2}\text{Mn}\text{Ir}_{5}\text{B}_{2}$, $\text{Mg}_{2}\text{Ni}\text{Ir}_{5}\text{B}_{2}$), whose easy magnetization axes are in the $z$ direction, have zero Berry curvature for the integral of the occupied states over the Brillouin zone, the anomalous Hall effect is not observed.
@@
Figure 4: The crystal and electronic structure of the altermagnet FeHO2 (31). a, The crystal primitive cell of the altermagnetic FeHO2 (31). be, Four significant magnetic structure of FeHO2 (31). The arrows represent the magnetic moments of Fe. f, The Brillouin zone (BZ) with high-symmetry points and lines of altermagnetic FeHO2 (31). g, The relative energy of four significant magnetic states with the variation of correlation interaction U. h and i, The electronic band structure of FeHO2 (31) without SOC. The red and blue lines represent the spin-up and spin-down energy bands, respectively. The electronic structure is calculated under correlation interaction $\rm{U=4~{}eV}$.
On the other hand, ferromagnetic semiconductors that have spintronic and transistor functionalities could be applied to the next generation of electronic devices. However, the ferromagnets usually are metals with a high Curie temperature and hold no brief for insulators with a high Curie temperature. Altermagnets with compensated antiparallel sublattices not only are in favor of insulators with a high Neel temperature but also have spintronic functionality [3]. Thus, altermagnets open a new pathway to bypass the difficulties of ferromagnets. Here, we employed the LDA+U method [84] to predict 34 altermagnetic semiconductors (see Supplementary Appendix Table S.2). Furthermore, the altermagnetic $\text{Fe}\text{H}\text{O}_{2}$ may be a spin-triplet excitonic phase. From Fig. 4i, we observe that there is large spin splitting of 0.39 eV on the $\rm{T-\Gamma-T}$ directions and the spins of valence and conduction band are opposite, which may result in the spin-triplet excitonic phase [85]. Moreover, the energy of altermagnetic state (AFM1) is much lower than that of the other three magnetic states (see Fig. 4g), indicating that $\text{Fe}\text{H}\text{O}_{2}$ may have a Neel temperature above the room temperature. Thus, the altermagnetic $\text{Fe}\text{H}\text{O}_{2}$ is a very interesting material that, we believe, will attract both theoretical and experimental interests. In addition, although $\text{Nd}\text{Ru}\text{O}_{3}$ is a altermagnetic semimetal, it has bandgap along the high-symmetry directions with spin splitting, and its valence and conduction bands have opposite spins (Supplementary Appendix Fig. S.5e and f). Thus, $\text{Nd}\text{Ru}\text{O}_{3}$ may be a Bardeen Cooper Schrieffer (BCS) type triplet exciton insulator [86]. In the following, we present in detail two altermagnetic materials which are metal and semiconductor, respectively.
Figure 5: The crystal and electronic structure of the altermagnetic $\text{Nb}_{2}\text{Fe}\text{B}_{2}$. a, The side view of altermagnetic $\text{Nb}_{2}\text{Fe}\text{B}_{2}$. b, The top view of altermagnetic $\text{Nb}_{2}\text{Fe}\text{B}_{2}$. c, The Brillouin zone (BZ) with high-symmetry points of altermagnetic $\text{Nb}_{2}\text{Fe}\text{B}_{2}$. The cyan plane represents the nodal surface of BZ. The m represents the mirror symmetry $M_{y}$. d, The anisotropic spin-charge density deriving from an anisotropic crystal field. e, The electronic band structure along high-symmetry directions without spin-orbit coupling (SOC). f, The electronic band structure along non-high-symmetry directions without SOC. The red and blue lines represent the spin-up and spin-down energy bands, respectively. The electronic structure is calculated under correlation interaction $\rm{U=4.82~{}eV}$.
@@
Figure 6: The crystal and electronic structure of the altermagnetic $\text{Na}\text{Fe}\text{O}_{2}~{}(92)$. a, The crystal structures of the $\text{Na}\text{Fe}\text{O}_{2}~{}(92)$ with FM structures. b, The crystal structures of the $\text{Na}\text{Fe}\text{O}_{2}~{}(92)$ with AFM1 structures. c, The crystal structures of the $\text{Na}\text{Fe}\text{O}_{2}~{}(92)$ with AFM2 structures. d, The crystal structures of the $\text{Na}\text{Fe}\text{O}_{2}~{}(92)$ with AFM3 structures. e, Relative energy of different magnetic states with the variation of correlation interaction $U$. f, The electronic band structure of $\text{Na}\text{Fe}\text{O}_{2}~{}(92)$ along high-symmetry directions without SOC. The red and blue lines represent the spin-up and spin-down energy bands, respectively. The electronic structure is calculated under correlation interaction $\rm{U=4~{}eV}$.
@@
The other altermagnetic material that we would like to mention is $\text{Na}\text{Fe}\text{O}_{2}$. The crystal structure of $\text{Na}\text{Fe}\text{O}_{2}$ is shown in Fig. 6a-d with space group $P/4_{1}2_{1}2~{}(92)$ symmetry. The corresponding elementary symmetry operations are $C_{4z}(\frac{1}{2},\frac{1}{2},\frac{1}{4})$ and $C_{2x}(\frac{1}{2},\frac{1}{2},\frac{3}{4})$, which yield the point group $D_{4}$. Since the $d$ orbitals of $Fe$ are half occupied and the angle between $Fe-O-Fe$ is 136 degrees in $\text{Na}\text{Fe}\text{O}_{2}$, the superexchange interactions result in the nearest neighbor $Fe$ ions with opposite magnetic moments and the next neighbor $Fe$ ions with the same magnetic moments. Hence, the magnetic ground state of $\text{Na}\text{Fe}\text{O}_{2}$ will be the AFM1 (see Fig. 6b). To verify our theoretical analysis, we consider four different magnetic structures, which are shown in Fig. 6a-d. It can be seen that the magnetic structure AFM1 is always in the ground state of $\text{Na}\text{Fe}\text{O}_{2}$ under different correlation interactions $U$ (see Fig. 6e). Moreover, the energy of the AFM1 state is much lower than that of the other three magnetic states (Fig. 6e), implying that $\text{Na}\text{Fe}\text{O}_{2}$ may have a Neel temperature above the room temperature. It is shown in Fig. 6b that the magnetic and crystal primitive cells of $\text{Na}\text{Fe}\text{O}_{2}$ are the same, which break $\{C_{2}^{\perp}||\tau\}$ spin symmetry. Thus, $\text{Na}\text{Fe}\text{O}_{2}$ is an altermagnetic material due to the lack of space-inversion symmetry.
@@
We also calculated the electronic band structure along the high-symmetry directions. Fig. 6f shows that the altermagnet $\text{Na}\text{Fe}\text{O}_{2}$ is a semiconductor with a band gap of 2.75 eV. The spin-degenerate bands on the $\Gamma-{\rm{X,~{}M}}^{*}-{\rm{A}}$ and $\rm{Z-R}$ directions (the $\rm{X-M}$ direction) are protected by the spin symmetry $\{C_{2}^{\perp}||C_{2y}(\frac{1}{2},\frac{1}{2},\frac{1}{4})\}$ ( $\{C_{2}^{\perp}||C_{2y}(\frac{1}{2},\frac{1}{2},\frac{3}{4})\}$) (see Fig. 6f). In fact, the spin symmetry $\{T||C_{2y}T(\frac{1}{2},\frac{1}{2},\frac{1}{4})\}$ ( $\{T||C_{2x}T(\frac{1}{2},\frac{1}{2},\frac{3}{4})\}$) can protect spin degeneracy of bands on the $k_{y}=0$ and $\pi$ (the $k_{x}=0$ and $\pi$) planes. That is to say, the altermagnet $\text{Na}\text{Fe}\text{O}_{2}$ has four nodal surfaces in the Brillouin Zone. Thus, $\text{Na}\text{Fe}\text{O}_{2}$ is a $d$-wave altermagnet which is reflected by the spin-splitting bands on the $\rm{M}-\Gamma-\rm{M}^{*}$ directions. Considering the $d$-wave altermagnets allowing unique spin current by electrical means [5], the altermagnet $\text{Na}\text{Fe}\text{O}_{2}$ may have both spintronic and transistor functionalities at the room temperature.
@@ Discussion
AI approaches have shown ground-breaking capabilities in the discovery of materials in a large search space. An intractable challenge faced by AI lies in the shortage of sufficient labels or positive samples, e.g., in the case of the discovery of altermagnetic materials. We herein introduced an AI search engine that combines pre-trained crystal models (GNN pre-training and optimal transport theory) and physics-based methods (symmetry analysis and first-principles electronic structure calculations) to discover new altermagnetic materials with specific properties under minimal labeled sample conditions. Among 91,649 possible candidates, we identified 50 new altermagnetic materials covering metal, semiconductor, and insulator. Meanwhile, the proposed AI search engine also has the few-shot learning ability. For example, it is capable of predicting 25 altermagnetic materials only based on 14 positive samples (see Supplementary Appendix Note C). We observed various novel physical properties in these newly discovered altermagnetic materials, e.g., anomalous Hall effect, anomalous Kerr effect, and topological property. It is noted that 4 out of these 50 altermagnetic materials are $i$-wave types, which are discovered for the first time filling a gap in the literature. We demonstrate that the AI search engine is capable of uncovering a set of altermagnetic materials with unique properties, highlighting its potential for accelerated discovery of the materials with targeting properties.
@@
There still remain some potential limitations associated with the AI search engine. Firstly, we have to admit that the issue of imbalance between positive and negative samples during the fine-tuning stage exists, primarily due to the scarcity of known positive samples. We have also discussed the potential intrinsic error and computational cost associated with this AI search engine (see Supplementary Appendix Note C). Utilizing the translational and rotational symmetries of crystals to augment positive sample data may help address this challenge, which will be demonstrated in our future work. Secondly, if the materials exhibit distinct magnetic phases at varying temperatures, such as Mn5Si3, whose low-temperature spin pattern is reported to be non-collinear and high-temperature spine pattern as antiferromagnetic  [87], our current model is unable to predict their magnetic properties. In such a case, if the material with temperature-driven magnetic transition is used as the positive sample to train the AI model, the prediction might involve possible bias. We will improve our classifier model by considering temperature as a conditional input parameter to enhance its capability in screening and predicting materials with temperature-driven properties in the future. Thirdly, in the calculations of magnetic ground states, we focus only on collinear magnetic structures and do not consider non-collinear ones. This is because most of the materials we calculate do not exhibit geometric frustration. Even for a few materials like CoF3, which exhibit triangular geometric frustration, previous studies did not consider non-collinear magnetic structures [55]. Another limitation is that we have not yet found ideal altermagnetic topological insulators and altermagnetic topological semimetals (such as odd-under-time-reversal Dirac, and sixfold semimetals). Employing the decoder based on the pre-trained model to generate potential altermagnetic materials holds promise in overcoming this challenge. Furthermore, adopting a multimodal pre-training approach offers the potential to further enhance the accuracy of model predictions. The current pre-training only considers the single modality of crystal structure information. Leveraging information from other modalities (such as textual descriptions of crystal structures) may enhance the performance of the pre-trained model. These methods will be further explored in our future research endeavors.
In addition, an alternative to alter the proposed AI model is to replace the classifier with a regressor to predict the magnetic structure of a given material candidate, where such a regressor can be fine-tuned based on 2,138 known magnetic structures in the MAGNDATA database [27]. Afterward, symmetry analysis can be employed to identify the altermagnetism. However, since there are infinite types of magnetic structures, accurately predicting the exact type of magnetic structure for a given material remains challenging. Therefore, we trust an end-to-end classifier in our proposed model, to directly judge whether a crystal material is altermagnetic, is preferred especially under the condition of very limited positive samples.
@@
It is well known in the community that the altermagnetism of 98% of candidates in the Material Project ($>$90,000 materials), whose magnetic structure information is unknown, has not yet been confirmed and remains a substantially challenging task. The brute-force approach leveraging our experts’ knowledge relies on trial and error by chance, having an extremely low probability of correctly discovering and confirming altermagnetic materials from the large database. However, our AI model narrows down the search space, predicts a list of highly possible candidate altermagnets, and lifts the discovery accuracy to a notable margin of about 31% (50/161), which greatly accelerates our discovery of new altermagnetic materials.  [^3] If we calculate the magnetic ground state of all materials, and then determine whether these materials are altermagnetic by spin symmetry, we probably can predict more altermagnetic materials. However, this cost is extremely huge. Given the scale and complexity of this task, it is unlikely to use DFT to complete the above brute-force computation for all candidate materials. The success of this engine lies not only in its predictive capabilities but also in its ability to leverage extensive crystal structure data and deep learning techniques, allowing for pre-training without explicit reference to underlying physical laws, to reveal complex correlations and patterns in new materials. Although 161 altermagnetic materials have been confirmed by the symmetry analysis approach [23], the urgency of discovering a variety of new altermagnetic materials with different properties still remains. Based on these existing altermagnetic materials confirmed by the spin space group, our AI search engine could predict many more altermagnetic materials, among which we expect to find a variety of altermagnets beyond the MAGNDATA database. For example, over 300 additional candidate materials (unconfirmed yet by DFT calculations) were predicted by AI search engine and listed in our GitHub repository [^4] https://github.com/zfgao66/MatAltMag for open research. Nevertheless, obtaining the magnetic ground state without experimental validation remains a challenging problem. The magnetic state of a material may be collinear or non-collinear, and the magnetic cell may be a supercell of the crystal primitive cell, which, in principle, lead to an infinite number of possibilities for the magnetic structure. Hence, experimental efforts will help further validate our discovery.
The proposed engine might be also applicable to other types of materials whose properties are strongly correlated to their crystal structures, such as the Bardeen-Cooper-Schrieffer (BCS) superconductive materials [88], ferromagnetic semi-conductor [89], and high-temperature superconductive materials [90], among others. We will demonstrate the potential of our proposed model for discovering other materials in our future study. We envision this effort may present new opportunities in the field of material discovery across different disciplines.